OpenStudy (anonymous):

can you help me how to solve this problem: compute the(a)Fg and FE(b)b/w an electron and proton separated by a distance of 0.04m..]

OpenStudy (anonymous):

Now I assume you mean the force of gravity (Fg) and the electrical force (FE). In that case you will need to use the following equations \[F_{g}=Gm_{1}m_{2}/r^{2}\] where G = the gravitational constant m1= mass of the electron m2 = mass of the proton r= the distance separating the electron and the proton For the electrical force you use \[F_{e}=kq_{1}q_{2}/r^{2}\] where \[k=1/4\pi\ \epsilon_{o}\approx 8.99 \times 10^{9} N.m^{2}/C^{2}\] q1 = the charge of the electron q2 = the charge of the electron and r = the distance separating the two charges

OpenStudy (anonymous):

\[G=6.673 \times 10^{-11} m^{3} kg^{-1} s^{-2}\]

OpenStudy (anonymous):

..tnx..hmm do u have the value of electron and proton?

OpenStudy (anonymous):

You should be able to find them in any physics books, either in the front or the back pages including all the other constants. You can also google them.

OpenStudy (anonymous):

can Fg=Gm1m2/r2..can also be like this.. Fg=Gm1m2/d2?

OpenStudy (anonymous):

Yep..The r in the equation I used is just a label. That is you can call it anything you like. :)

OpenStudy (anonymous):

How about you calculate the two forces and show me your answer and I'll check if they are correct?

OpenStudy (anonymous):

this is for a.. im not sure if the value of electron and proton is correct..electron=9.108x10raise to negative 28 g..and proton is 1.672x10 raise to negative 24g.. fg=Gm1m2/d2 fg=6.67x10raise to neg.11 nm2/kg(9.11x10raise to neg.31 kg)(1.67x10raise to neg.27 kg)/(o.o4m)2 fg=1.015x10raise to neg.67Nm2/1.6x10raise to neg.3m2 =6.34x10raise to -11N.. HMM??i dont know if it is correct..will you check it?

OpenStudy (anonymous):

The electron mass = 9.11 × 10^(-31) kilograms. The proton mass = 1.67 × 10^(-27) kilograms. Try finding the answer to this equation\[F_{g} =(6.67 \times 10^{-11} \times 1.67 \times 10^{-27} \times 9.11 \times 10^{-31})/ 0.04^{2}\] and tell me what you get. Take your time. :)

OpenStudy (anonymous):

and check that this is the same as what you calculated.

OpenStudy (anonymous):

..im sorry..i dont have a scientific calculator?..i left it in school..

OpenStudy (anonymous):

it is best that you do most of it by hand. That is work out what all the power of tens come to first. What is \[?=10^{-11} \times 10^{-27} \times 10^{-31} \]

OpenStudy (anonymous):

sorry..i cant answer it w/out calculator..im not really good in physics

OpenStudy (anonymous):

The good thing about this is that it's kinda easy to work out as long as you realise that \[10^{-11} \times 10^{-27} \times 10^{-31} = 10^{-11-27-31}=?\]

OpenStudy (anonymous):

thanks for the information..hmm can you give me the idea how to answer the b?

OpenStudy (anonymous):

By using the equation I gave for the electrical force. i.e. The second equation I wrote.

OpenStudy (anonymous):

..thank you so much..very big help for me..i can have a better future now..hehe..thanks:)